# Information About Other Agents in Mechanism Design

### v3

@@ 4.2 Bayesian monotonicity
$\mathbb{E}(u_{1}(F(\theta_{-1},\theta_{1}),\theta_{1})\mid\theta_{1})\geq\mathbb{E}(u_{1}(y(\theta_{-1}),\theta_{1})\mid\theta_{1})\Rightarrow A+2B\leq 20.$
From here it is easy to see that there are no solutions to both inequalities satisfying $A\leq 9.5$ and $B\geq 5$. By similar reasoning, there is no $y:\Theta_{-1}\to A$ such that
$\mathbb{E}(u_{1}(y(\alpha_{-1}(\theta_{-1})),\theta_{1}^{\prime})\mid\theta_{1}^{\prime})>\mathbb{E}(u_{1}(F(\alpha(\theta_{-1},\theta_{1}^{\prime})),\theta_{1}^{\prime})\mid\theta_{1}^{\prime})$
and
$\mathbb{E}(u_{1}(F(\theta_{-1},\theta_{1}^{\prime}),\theta_{1}^{\prime})\mid\theta_{1}^{\prime})<\mathbb{E}(u_{1}(y(\theta_{-1}),\theta_{1}^{\prime})\mid\theta_{1}^{\prime}).$
Thus, agent 1 cannot undermine $\alpha$. For agent 2, the utility experienced from outcome $f$ is too high for her to include it as a possible outcome in any alternate SCF $y.$ Thus, similar reasoning as what was provided for agent 1 shows that agent 2 also cannot undermine $\alpha$. For agent 3, note that her expected utility for truth-telling versus the deception $\alpha$ will always be the same, no matter what the SCF is. Therefore, the expected utility for agent 3 in the specified SCF is the same for truth-telling versus the deception $\alpha$, and similarly for any alternate SCF $y$. It follows that agent 3 can’t profit from signaling the deception via an alternate SCF unless she profits from the alternate SCF without deception as well, so agnet 3 cannot undermine $\alpha$.  [^7] This property may appear to render this example a knife-edge case, but we show in Appendix A that this example can be modified in a straightforward way to remove its reliance on agent 3’s indifference. For now, agent 3’s indifference allows the example to be somewhat simpler and focus on the effect of an additional signal. Therefore $F$ is not fully implementable if no agents receive informative signals about the types of other agents.
4.3 Eliminating $\alpha$ as an equilibrium
We claim that there is a binary signal that agent 3 can receive regarding the types of agents 1 and 2 that allows $F$ to satisfy Bayesian monotonicity and as result become fully implementable. Suppose $\sigma_{3}\in\{0,1\}$ is such that
$P(\sigma_{3}=0\mid\theta_{-3}=(\theta_{1},\theta_{2}))=P(\sigma_{3}=0\mid\theta_{-3}=(\theta_{1}^{\prime},\theta_{2}))=2/3$
and
$P(\sigma_{3}=1\mid\theta_{-3}=(\theta_{1},\theta_{2}^{\prime}))=P(\sigma_{3}=1\mid\theta_{-3}=(\theta_{1}^{\prime},\theta_{2}^{\prime}))=2/3.$
If $\sigma_{3}=0$, then agent 3’s updated beliefs place $2/9$ probability on both $(\theta_{1},\theta_{2})$ and $(\theta_{1},\theta_{2}^{\prime})$, $1/9$ probability on $(\theta_{1}^{\prime},\theta_{2}^{\prime})$, and $4/9$ probability on $(\theta_{1}^{\prime},\theta_{2}).$ If $\sigma_{3}=1,$ then agent 3’s updated beliefs place $1/9$ probability on $(\theta_{1},\theta_{2}),$ $2/9$ probability on both $(\theta_{1}^{\prime},\theta_{2})$ and $(\theta_{1}^{\prime},\theta_{2}^{\prime})$, and $4/9$ probability on $(\theta_{1},\theta_{2}^{\prime}).$
Since showing that none of the $\alpha_{i}$ for $1\leq i\leq 5$ violate Bayesian monotonicity did not require any information about agent 3’s beliefs, and the signal agent 3 receives modifies only her beliefs, it suffices to consider only the deviation $\alpha$ when showing that $F$ satisfies Bayesian monotonicity. Note that $F$ needs to satisfy Bayesian monotonicity in either realization of $\sigma_{3}.$
First suppose $\sigma_{3}=0$. Define $y:\Theta_{-3}\to A$ by $y(\theta_{1},\theta_{2})=d,y(\theta_{1},\theta_{2}^{\prime})=b,y(\theta_{1}^{\prime},\theta_{2})=e,$ and $y(\theta_{1}^{\prime},\theta_{2}^{\prime})=a.$ Then
$\mathbb{E}(u_{3}(F(\theta_{-3},\theta_{3}),\theta_{3})\mid\theta_{3})=\frac{2}{9}u_{3}(a,\theta_{3})+\frac{2}{9}u_{3}(b,\theta_{3})+\frac{4}{9}u_{3}(e,\theta_{3})+\frac{1}{9}u_{3}(d,\theta_{3})=32/9$
$\geq\mathbb{E}(u_{3}(y(\theta_{-3}),\theta_{3})\mid\theta_{3})=\frac{1}{9}u_{3}(a,\theta_{3})+\frac{2}{9}u_{3}(b,\theta_{3})+\frac{4}{9}u_{3}(e,\theta_{3})+\frac{2}{9}u_{3}(d,\theta_{3})=29/9$
and
$\mathbb{E}(u_{3}(F(\theta_{-3},\theta_{3}^{\prime}),\theta_{3}^{\prime})\mid\theta_{3}^{\prime})=32/9\geq\mathbb{E}(u_{3}(y(\theta_{-3}),\theta_{3}^{\prime})\mid\theta_{3}^{\prime})=23/9$
but
$\mathbb{E}(u_{3}(F(\alpha(\theta_{-3},\theta_{3})),\theta_{3})\mid\theta_{3})=\frac{1}{9}u_{3}(a,\theta_{3})+\frac{4}{9}u_{3}(b,\theta_{3})+\frac{2}{9}u_{3}(e,\theta_{3})+\frac{2}{9}u_{3}(d,\theta_{3})=31/9$
$<\mathbb{E}(u_{3}(y(\alpha(\theta_{-3})),\theta_{3})\mid\theta_{3})=\frac{2}{9}u_{3}(a,\theta_{3})+\frac{4}{9}u_{3}(b,\theta_{3})+\frac{2}{9}u_{3}(e,\theta_{3})+\frac{1}{9}u_{3}(d,\theta_{3})=34/9.$
Now suppose $\sigma_{3}=1.$ Define $y:\Theta_{-3}\to A$ by $y(\theta_{1},\theta_{2})=a,y(\theta_{1},\theta_{2}^{\prime})=e,y(\theta_{1}^{\prime},\theta_{2})=b,y(\theta_{1}^{\prime},\theta_{2}^{\prime})=d.$ Then
$\mathbb{E}(u_{3}(F(\theta_{-3},\theta_{3}),\theta_{3})\mid\theta_{3})=\frac{1}{9}u_{3}(a,\theta_{3})+\frac{4}{9}u_{3}(b,\theta_{3})+\frac{2}{9}u_{3}(e,\theta_{3})+\frac{2}{9}u_{3}(d,\theta_{3})=31/9$
$\geq\mathbb{E}(u_{3}(y(\theta_{-3},\theta_{3}),\theta_{3})\mid\theta_{3})=\frac{1}{9}u_{3}(a,\theta_{3})+\frac{2}{9}u_{3}(b,\theta_{3})+\frac{4}{9}u_{3}(e,\theta_{3})+\frac{2}{9}u_{3}(d,\theta_{3})=29/9$
and
$\mathbb{E}(u_{3}(F(\theta_{-3},\theta_{3}^{\prime}),\theta_{3}^{\prime})\mid\theta_{3}^{\prime})=31/9\geq\mathbb{E}(u_{3}(y(\theta_{-3},\theta_{3}^{\prime}),\theta_{3}^{\prime})\mid\theta_{3}^{\prime})=23/9.$
But
$\mathbb{E}(u_{3}(F(\alpha(\theta_{-3},\theta_{3}),\theta_{3}))\mid\theta_{3})=\frac{2}{9}u_{3}(a,\theta_{3})+\frac{2}{9}u_{3}(b,\theta_{3})+\frac{4}{9}u_{3}(e,\theta_{3})+\frac{1}{9}u_{3}(d,\theta_{3})=32/9$
$<\mathbb{E}(u_{3}(y(\alpha(\theta_{-3},\theta_{3}),\theta_{3}))\mid\theta_{3})=\frac{2}{9}u_{3}(a,\theta_{3})+\frac{4}{9}u_{3}(b,\theta_{3})+\frac{2}{9}u_{3}(e,\theta_{3})+\frac{1}{9}u_{3}(d,\theta_{3})=34/9.$
Thus, upon receiving the signal $\sigma_{3}$, agent 3 can undermine $\alpha$. The intuition of the signal construction is as follows. For $F$ as described, truth-telling is a dominant strategy for agent 3. Additionally, agent 3 always prefers that agent 1 and 2’s type profile is $(\theta_{1},\theta_{2})$ rather than $(\theta_{1}^{\prime},\theta_{2}^{\prime})$, regardless of her type. Suppose agent 3 is quite confident that agent 2 is type $\theta_{2},$ so she believes it is more likely that the type profile of agents 1 and 2 belongs to $\{(\theta_{1},\theta_{2}),(\theta_{1}^{\prime},\theta_{2})\}$ rather than $\{(\theta_{1},\theta_{2}^{\prime}),(\theta_{1}^{\prime},\theta_{2}^{\prime})\}$. In particular, suppose she is confident enough that she places a higher probability on $(\theta_{1},\theta_{2})$ than $(\theta_{1}^{\prime},\theta_{2}^{\prime}).$ Then, under the SCF $F$, if the deception $\alpha$ is being played, she believes it is more likely that her utility will be harmed by agents $(\theta_{1},\theta_{2})$ deviating to $(\theta_{1}^{\prime},\theta_{2}^{\prime}),$ than benefited by the vice versa happening. It is therefore beneficial for her to propose the alternate SCF $y:\Theta_{-3}\to A$, which is defined by $y(\theta_{1},\theta_{2})=d,y(\theta_{1},\theta_{2}^{\prime})=b,y(\theta_{1}^{\prime},\theta_{2})=e,$ and $y(\theta_{1}^{\prime},\theta_{2}^{\prime})=a$, that swaps the outcomes for the type profiles $(\theta_{1},\theta_{2})$ and $(\theta_{1}^{\prime},\theta_{2}^{\prime})$ while holding the outcomes for the other type profiles constant. Similarly, agent 3 prefers that agent 1 and 2’s type profile is $(\theta_{1},\theta_{2}^{\prime})$ rather than $(\theta_{1}^{\prime},\theta_{2}).$ If agent 3 is quite confident that the type profile of agents 1 and 2 belongs to $\{(\theta_{1},\theta_{2}^{\prime}),(\theta_{1}^{\prime},\theta_{2}^{\prime})\}$, then if the deception $\alpha$ is being played with the SCF $F$, she will believe it is more likely that her utility will be harmed by agents $(\theta_{1},\theta_{2}^{\prime})$ deviating to $(\theta_{1}^{\prime},\theta_{2})$ than benefited by the vice versa happening. Thus, swapping the outcomes in these cases by proposing the alternate SCF $y:\Theta_{-3}\to A$ defined by $y(\theta_{1},\theta_{2})=a,y(\theta_{1},\theta_{2}^{\prime})=e,y(\theta_{1}^{\prime},\theta_{2})=b,$ and $y(\theta_{1}^{\prime},\theta_{2}^{\prime})=d$ allows her undermine $\alpha.$
4.4 Full implementability of $F$
We want to conclude that $F$ is implementable with respect to $\{\sigma_{i}\}_{i\in\{1,2,3\}}$ where $\sigma_{1},\sigma_{2}$ are uninformative and $\sigma_{3}$ is as specified above. To do so, we will analyze the equilibria of the direct mechanism for $F$ under this modified information structure for agent 3 and apply lemma 1.
It was shown above that truth-telling and $\alpha$ were equilibria of the direct mechanism under the uninformative information structure. It is straightforward to verify that these equilibria are unchanged in either realization of $\sigma_{3}$. The other equilibria of the direct mechanism, in either realization of $\sigma_{3}$, are the following: $\alpha_{1}=((\theta_{1}^{\prime},\theta_{1}^{\prime}),(\theta_{2},\theta_{2}),(\theta_{3},\theta_{3}^{\prime})),\alpha_{2}=((\theta_{1},\theta_{1}),(\theta_{2}^{\prime},\theta_{2}^{\prime}),(\theta_{3},\theta_{3}^{\prime})),\alpha_{3}=((\theta_{1}^{\prime},\theta_{1}^{\prime}),(\theta_{2},\theta_{2}),(\theta_{3}^{\prime},\theta_{3}^{\prime})),\alpha_{4}=((\theta_{1}^{\prime},\theta_{1}^{\prime}),(\theta_{2},\theta_{2}),(\theta_{3}^{\prime},\theta_{3}))$, and $\alpha_{5}=((\theta_{1}^{\prime},\theta_{1}^{\prime}),(\theta_{2},\theta_{2}),(\theta_{3},\theta_{3}))$.
To see this, suppose agent 1 deviates to $(\theta_{1}^{\prime},\theta_{1}^{\prime}).$ Since both types of agent 2 prefer $e$ to $d$, agent 2’s best response is $(\theta_{2},\theta_{2}).$ Both types of agent 1 prefer $e$ to $a$, so $(\theta_{1}^{\prime},\theta_{1}^{\prime})$ is also a best response to $(\theta_{2},\theta_{2}).$ Thus $\alpha_{1}=((\theta_{1}^{\prime},\theta_{1}^{\prime}),(\theta_{2},\theta_{2}),(\theta_{3},\theta_{3}^{\prime}))$ is an undesired equilibrium in the direct mechanism. Similarly, if agent 1 deviates to $(\theta_{1},\theta_{1}),$ both types of agent 2 prefer $b$ or $c$ to $a,$ so agent 2’s best response is $(\theta_{2}^{\prime},\theta_{2}^{\prime}).$ Since both types of agent 1 prefer $b$ or $c$ to $d,$ agent 1’s best response is $(\theta_{1},\theta_{1}),$ so $\alpha_{2}=((\theta_{1},\theta_{1}),(\theta_{2}^{\prime},\theta_{2}^{\prime}),(\theta_{3},\theta_{3}^{\prime}))$ is another undesired equilibrium in the direct mechanism. Finally, since agent 3’s report only affects the outcome of the direct mechanism if agent 1 reports $\theta_{1}$ and agent 2 reports $\theta_{2}^{\prime}$, it follows that $\alpha_{3}=((\theta_{1}^{\prime},\theta_{1}^{\prime}),(\theta_{2},\theta_{2}),(\theta_{3}^{\prime},\theta_{3}^{\prime})),\alpha_{4}=((\theta_{1}^{\prime},\theta_{1}^{\prime}),(\theta_{2},\theta_{2}),(\theta_{3}^{\prime},\theta_{3}))$, and $\alpha_{5}=((\theta_{1}^{\prime},\theta_{1}^{\prime}),(\theta_{2},\theta_{2}),(\theta_{3},\theta_{3}))$ are all undesired equilibria, in either realization of $\sigma_{3}$. Truth-telling, $\alpha,\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4},$ and $\alpha_{5}$ cover all possible strategies for agent 1 in the direct mechanism. Since they are all equilibria and truth-telling is a dominant strategy for agent 3 (a property that is independent of $\sigma_{3}$’s realization), there can be no other equilibria in the direct mechanism, in either realization of $\sigma_{3}$.
We claim that each deception $\alpha_{i}$ for $1\leq i\leq 5$ can be undermined by either agent 1 or agent 2. To see this, first note that all such deceptions make the expected utility of agents 1 and 2 certain, since agents 1 and 2 are not sending different messages for different realizations of their own type. Moreover, in no case does a deception $\alpha_{i}$ allow all types of agents 1 and 2 to for sure get their most preferred outcome. This makes it easy for the agents to define an alternate SCF that would benefit them. For example, in the case of $\alpha_{1},$ consider the function $y:\Theta_{-1}\to A$ defined by $y(\theta_{2},\cdot)=d,y(\theta_{2}^{\prime},\cdot)=f$. Then
$E(u_{1}(F(\theta_{-1},\theta_{1}),\theta_{1})\mid\theta_{1})=10/3\geq E(u_{1}(y(\theta_{-1}),\theta_{1})\mid\theta_{1})=4/3$
and
$E(u_{1}(F(\theta_{-1},\theta_{1}^{\prime}),\theta_{1}^{\prime})\mid\theta_{1}^{\prime})=10/3\geq E(u_{1}(y(\theta_{-1}),\theta_{1}^{\prime})\mid\theta_{1}^{\prime})=3$
but
$E(u_{1}(y(\alpha_{1}(\theta_{-1})),\theta_{1}^{\prime})\mid\theta_{1}^{\prime})=4>E(u_{1}(F(\alpha_{1}(\theta_{-1},\theta_{1}^{\prime})),\theta_{1}^{\prime})\mid\theta_{1}^{\prime})=3.$
It follows that $\alpha_{1}$ is undermined by agent 1. Similar arguments show that $\alpha_{2},\alpha_{3},\alpha_{4},$ and $\alpha_{5}$ are also undermined. Since $\sigma_{3}$ only affects the information of agent 3, each $\alpha_{i}$ for $1\leq i\leq 5$ can be undermined by either agent 1 or agent 2 in both realizations of $\sigma_{3}$. In the previous subsection we saw that agent 3 can undermine $\alpha$ in either realization. Therefore, lemma 1 implies that $F$ is fully implementable with respect to $\{\sigma_{i}\}_{i\in\{1,2,3\}}$.
This example also illustrates how a planner can strictly prefer agents to have an information structure that is between no information and complete information. Indeed, $F$ is also not fully implementable in a complete information environment because it fails Maskin monotonicity: when moving from the state $(\theta_{1}^{\prime},\theta_{2},\theta_{3})$ to $(\theta_{1},\theta_{2},\theta_{3})$, the outcome $e$ is chosen in the former state and weakly rises in every agent’s preference ordering, but $e$ is not chosen in the latter state.
5 General results
@@
The main example suggests how an agent for which truth-telling is a dominant strategy for a given SCF $F$ can undermine a given deception $\alpha$ with additional information about the types of other agents, in a way that she would not be able to do so without this information. In particular, suppose $F$ is an SCF such that there is a type profile $\theta_{-i}^{\prime}$ that yields an outcome agent $i$ prefers to that of the type profiles in $\alpha_{-i}(\theta_{-i}^{\prime})$, regardless of what agent $i$ reports. Suppose additionally that agent $i$ believes $\theta_{-i}^{\prime}$ is more likely to be the true type profile of other agents than the type profiles in $\alpha_{-i}^{-1}(\theta_{-i}^{\prime})$. Then it is possible for some type $\theta_{i}$ of agent $i$ to undermine $\alpha$ by proposing the alternate SCF that swaps the outcomes corresponding to $\theta_{-i}^{\prime}$ and $\alpha_{-i}(\theta_{-i}^{\prime})$ assuming she plays $\alpha_{i}(\theta_{i})$. Thus, if agent $i$ can receive a signal that ensures such a type profile $\theta_{-i}^{\prime}$ always exists, then an undesirable equilibrium $\alpha$ can be potentially undermined. We formalized this into the following condition.
@@
hold for all $\theta_{i}^{\prime}\in\Theta_{i}.$ Furthermore the inequalities are strict for some type $\theta_{i}^{\prime}=\theta_{i}.$
The inequalities capture the intuition that agent $i$ should always prefer other agents report truthfully rather than according to the deception $\alpha$ if she knew the type profiles $\theta_{-i}^{\prime}$ or $\theta_{-i}^{\prime\prime}$ realized. This preference needs to be strict for at least one type of agent $i$ in order to provide sufficient incentives for her to deviate from the undesirable equilibrium $\alpha.$ The requirements that $\alpha_{-i}(\alpha_{-i}(\theta_{-i}^{\prime}))\neq\theta_{-i}^{\prime\prime},\alpha_{-i}(\alpha_{-i}(\theta_{-i}^{\prime\prime}))\neq\theta_{-i}^{\prime},$ and $\alpha_{-i}(\theta_{-i}^{\prime})\neq\alpha_{-i}(\theta_{-i}^{\prime\prime}),$ limit the interactions $\theta_{-i}^{\prime}$ and $\theta_{-i}^{\prime\prime}$ have with each other in the deception $\alpha$, which allows us to partition the type space $\Theta_{-i}$ in a natural way. As the name suggests, an agent $i$ that can potentially undermine a deception $\alpha$ can undermine $\alpha$ with additional information about the types of other agents. In other words, we have the following lemma, which is the main technical contribution of this paper.
@@
The full proof is given in Appendix B. For an incentive-compatible $F$ that is not fully implementable, the lemma suggests a systematic way to eliminate all deceptions in $\mathcal{A}_{F}\setminus\mathcal{A}_{F}^{u}$ (which must be non-empty by lemma 1) with additional information. In particular, we have the following result.
@@
The result is essentially a straightforward application of lemmas 1 and 2, and the proof is in the appendix. The theorem requires a separation of the agents who can undermine deceptions in $\mathcal{A}_{F}^{u}$ from the agents who can potentially undermine deceptions in $\mathcal{A}_{F}\setminus\mathcal{A}_{F}^{u}$ because we cannot guarantee that an agent who undermines a deception $\alpha\in\mathcal{A}_{F}^{u}$ without additional information can still undermine $\alpha$ after receiving an informative signal about other agents. This implicitly requires $|\mathcal{A}_{F}\setminus\mathcal{A}_{F}^{u}|\leq n-1$.

### v4

@@ 4 An illustrative example
$Pr(\sigma_{3}=1\mid\theta_{-3}=(H_{1},L_{2}))=Pr(\sigma_{3}=1\mid\theta_{-3}=(L_{1},L_{2}))=\tau$
for any $\tau>1/2$.
The signal $\sigma_{3}$ essentially provides a noisy signal to the CEO regarding whether the junior analyst believes the demand is high or low. This has many possible interpretations, such as allowing the CEO to monitor the junior analyst’s work more closely.
Proof.
Suppose $\sigma_{3}=0$ is realized. Then the CEO believes it is more likely that the type profile of the analysts belongs to $\{(H_{1},H_{2}),(L_{1},H_{2})\}$ rather than $\{(H_{1},L_{2}),(L_{1},L_{2})\}$. Since $\tau>1/2,$ she believes it is more likely for the analysts’ type profile to be $(H_{1},H_{2})$ than $(L_{1},L_{2}).$ Suppose she proposes the alternate SCF
$y(\theta_{1},\theta_{2})=\begin{cases}(E,1,1)&\text{if }(\theta_{1},\theta_{2})=(H_{1},H_{2}),\\[4.0pt] (S,0,0)&\text{if }(\theta_{1},\theta_{2})=(H_{1},L_{2}),\\[4.0pt] (E,0,0)&\text{if }(\theta_{1},\theta_{2})=(L_{1},H_{2}),\\[4.0pt] (I,1,1)&\text{if }(\theta_{1},\theta_{2})=(L_{1},L_{2}).\end{cases}$
which swaps the outcomes corresponding to $(H_{1},H_{2})$ and $(L_{1},L_{2}).$ Compared to $F$, the SCF $y$ gives her a lower utility if analysts are truth-telling but improves her utility if analysts are always lying. Thus the CEO can undermine always-lie if $\sigma_{3}=0$ is realized. Similarly, if $\sigma_{3}=1$ is realized, the CEO believes $(H_{1},L_{2})$ is a more likely analyst profile than $(L_{1},H_{2})$. A similar alternate $y$ swapping the outcomes corresponding to the two profiles allows her to undermine always-lie. Therefore the CEO can undermine always-lie in either realization of $\sigma_{3}.$ Since $\sigma_{3}$ only changes the CEO’s information, the deceptions always-$H$ and always-$L$ can still be undermined by the compliance officer in either realization. It follows that $F$ is fully implementable with respect to $\{\sigma_{i}\}.$
@@
This example illustrates how an agent for which truth-telling is a dominant strategy can undermine deceptions with additional information about the types of other agents. When an SCF $F$ fails to be fully implementable, the culprit is at least one equilibrium $\alpha$ of the direct mechanism that the planner dislikes and cannot be undermined. Suppose truth-telling is a dominant strategy for agent $i$, and she always prefers the outcome associated with the type profile $\theta_{-i}^{\prime}$ to the type profiles in $\alpha_{-i}(\theta_{-i}^{\prime})$ — regardless of her own type and what she reports. Then a simple “outcome” swap can kill $\alpha$ whenever she believes $\theta_{-i}^{\prime}$ is more likely to be the true type profile of other agents than the type profiles in $\alpha_{-i}^{-1}(\theta_{-i}^{\prime})$. A suitably constructed signal can ensure she holds such a belief. Additionally, as long as agent $i$ can receive a signal that ensures such a type profile $\theta_{-i}^{\prime}$ exists in each realization, then an undesirable equilibrium $\alpha$ can be killed.
5 A sufficient condition
We begin by formalizing the intuition provided in the previous section into the following condition.
@@ 5 A sufficient condition
hold for all $\theta_{i}^{\prime}\in\Theta_{i}.$ Furthermore, for each inequality there is a type $\theta_{i}$ that makes it strict.
In words, the definition requires two type profiles of other agents that agent $i$ prefers to be truthfully reported rather than manipulated by $\alpha$ in all states of the world, and for each type profile this preference is strict for at least one of agent $i$’s types. We note that the preferences do not need to be strict for the same type $\theta_{i}$. The requirements that
$\alpha_{-i}(\alpha_{-i}(\theta_{-i}^{\prime}))\neq\theta_{-i}^{\prime\prime},\alpha_{-i}(\alpha_{-i}(\theta_{-i}^{\prime\prime}))\neq\theta_{-i}^{\prime},\text{ and }\alpha_{-i}(\theta_{-i}^{\prime})\neq\alpha_{-i}(\theta_{-i}^{\prime\prime})$
limit the interactions $\theta_{-i}^{\prime}$ and $\theta_{-i}^{\prime\prime}$ have with each other in the deception $\alpha$, which allows for a natural partition of $\Theta_{-i}$. As the name suggests, an agent $i$ that can potentially undermine a deception $\alpha$ can undermine $\alpha$ with additional information about the types of other agents.
@@
We emphasize that the signal does not need to have a specific level of accuracy: Any sufficiently accurate signal will accomplish the same objective. For an incentive-compatible $F$ that is not fully implementable, the lemma suggests a systematic way to eliminate all deceptions in $\mathcal{A}_{F}\setminus\mathcal{A}_{F}^{u}$ (which must be non-empty by lemma 1) with additional information. In particular, we have the following result.
@@
Establishing the result hinges on showing that the set of Bayes Nash equilibria of the direct mechanism remain unchanged after agents receive the signals constructed and applying lemmas 1 and 2. In particular, the signals do not create additional undesirable equilibria of the direct mechanism that were not originally present. The theorem requires a separation of the agents who can undermine deceptions in $\mathcal{A}_{F}^{u}$ from the agents who can potentially undermine deceptions in $\mathcal{A}_{F}\setminus\mathcal{A}_{F}^{u}$ because it is not possible in general to guarantee that an agent who undermines a deception $\alpha\in\mathcal{A}_{F}^{u}$ without additional information can still undermine $\alpha$ after receiving an informative signal about other agents. This implicitly requires $|\mathcal{A}_{F}\setminus\mathcal{A}_{F}^{u}|\leq n-1$.